# REG4Rec: Reasoning-Enhanced Generative Model for Large-Scale Recommendation Systems

### v3

@@ 3.5. Inference with Reflection
We propose Consistency‑Oriented Self‑Reflection Pruning (CORP), a self-reflection mechanism applied at inference to identify and correct inconsistent reasoning paths. At every $s$-th step, the category predictor—trained offline—produces a probability distribution $\tilde{p}_{i}$ over item categories, serving as a semantic signal for consistency checking. Specifically, consistency is measured via JS divergence between consecutive distributions (Eq. (13)). If $JS(\tilde{p}_{i}\|\tilde{p}_{i+s})$ exceeds the self-reflection threshold $\theta$, it signals potential inconsistency or low confidence. The model then triggers self-reflection, terminating the current path and rolling back to an earlier step for regeneration. This pruning preserves only semantically coherent and logically consistent paths, enhancing overall recommendation accuracy.
4. Offline Experiments
@@
Table 1. Statistics of Public and Industrial Datasets. #Users and #Items denote the number of users and items, respectively; #Interactions denotes the total number of user–item interactions in the dataset.
@@
4.1. Experimental Setup
Dataset. Four datasets are used to demonstrate the performance of our proposed method. The first three are public datasets, while the fourth one is an in-house industrial dataset. Table 1 summarizes dataset statistics.
  • Amazon Product Reviews. We use three publicly available benchmarks from the Amazon Product Reviews dataset (He and McAuley, 2016), which contains user reviews and item meta information from May 1996 to July 2014. For GR, we focus on three categories: ”Beauty”, ”Sports and Outdoors”, and ”Toys and Games”. We generate training sequences from users’ review histories, sorted by timestamp, and exclude users with fewer than five reviews to ensure sufficient interactions. We follow the standard leave-one-out evaluation protocol (Kang and McAuley, 2018): the last item is used for testing, the second-to-last for validation, and all preceding items for training.
  • Industrial Dataset. We construct the in-house offline dataset from sequential user behaviors and feedback logs collected on a widely used e-commerce advertising platform. The dataset spans diverse recommendation scenarios, including homepage recommendations, product detail pages, and interactive gaming sections, capturing varied user behaviors. It contains data from 30 million users and 40 million advertisements collected between October 2024 and May 2025, offering a comprehensive view of real-world user behavior and ad content.
Table 2. Performance Comparison across baselines. The best metric for each dataset is highlighted in bold, while the second-best is underlined. The last row (Improv.) denotes the relative improvement of our REG4Rec over the best baseline.
Method Beauty Sports Toys Industrial Dataset
R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10
SASRec 0.0387 0.0249 0.0605 0.0318 0.0233 0.0154 0.0350 0.0192 0.0463 0.0306 0.0675 0.0374 0.0658 0.0453 0.0916 0.0564
$S^{3}$-Rec 0.0387 0.0244 0.0647 0.0327 0.0251 0.0161 0.0385 0.0204 0.0443 0.0294 0.0700 0.0376 0.0759 0.0530 0.1066 0.0628
TIGER 0.0454 0.0321 0.0648 0.0384 0.0264 0.0181 0.0400 0.0225 0.0521 0.0371 0.0712 0.0432 0.0903 0.0594 0.1293 0.0715
COBRA 0.0537 0.0395 0.0725 0.0456 0.0305 0.0215 0.0434 0.0257 0.0619 0.0462 0.0781 0.0515 0.0963 0.0646 0.1397 0.0785
ReaRec 0.0431 0.0276 0.0654 0.0375 0.0256 0.0179 0.0393 0.0219 0.0552 0.0374 0.0747 0.0435 0.0742 0.0517 0.0982 0.0603
STREAM 0.0538 0.0388 0.0691 0.0422 0.0324 0.0232 0.0452 0.0271 0.0644 0.0442 0.0799 0.0502 0.0981 0.0680 0.1422 0.0804
REG4Rec 0.0586 0.0416 0.0792 0.0493 0.0372 0.0252 0.0527 0.0296 0.0676 0.0493 0.0862 0.0574 0.1094 0.0738 0.1569 0.0890
Improv. +8.92% +5.32% +9.24% +8.11% +14.81% +8.62% +16.59% +9.23% +4.97% +6.71% +7.88% +11.46% +11.52% +8.53% +10.34% +10.70%
@@ 4.1. Experimental Setup
  • SASRec (Kang and McAuley, 2018) : A transformer-based model that employs the self-attention mechanism to capture long-term dependencies.
  • $S^{3}$-Rec (Zhou et al., 2020) : A transformer-based model that integrates self-supervised learning to enhance representation learning.
  • TIGER (Rajput et al., 2023) : A GR model built on an encoder-decoder structure, leveraging semantic IDs (based on RQ-VAE) and autoregressive decoder to improve the generalizable of recommended items.
  • COBRA (Yang et al., 2025) : A GR model with a decoder-only structure, incorporating semantic IDs (based on RQ-VAE) and dense vectors to capture semantic insights and collaborative signals.
  • ReaRec (Tang et al., 2025) : A transformer-based model that extends SASRec to generate multiple implicit intermediate outcomes without explicitly defined inference paths.
  • STREAM (Zhang et al., 2025b) : A transformer-based model uses textually similar items as explicit reasoning steps, guided by the GRPO algorithm for reasoning.
Evaluation Metrics. For evaluation, we adopt the Recall@K (R@K) and NDCG@K (N@K) as standard metrics for recommendation, following (Yang et al., 2025), with $K$ set to 5 and 10. These metrics assess both the accuracy of item retrieval and the quality of the ranking.
Implementation Details. This section summarizes our implementation settings. The training process is executed on a distributed TensorFlow (Abadi et al., 2016) platform, consisting of 10 parameter servers and 10 workers, each with 2 GPUs. To ensure a fair comparison with baselines, we configure the total number of transformer layers to 3, with one layer for the encoder and two layers for the decoder. The input dimension is 128 and the hidden size per transformer layer is 640. For MPQ, the hyperparameter $\alpha$ used in the training process of MPQ is set to 0.001. Because MPQ training is not the core focus and these weights have minimal impact once MPQ converges, we do not perform a sensitivity study on them. For flexible reasoning paths, MPQ uses 8 codebooks with 300 codewords each. The tokenizers process both text and image features for the industrial and public datasets. Text embeddings are generated using the pre-trained Qwen3-Embedding 7B model (Zhang et al., 2025a) and image embeddings are generated using the pre-trained Pailitao v8 model (Zhang et al., 2018), both with a dimensionality of 256. We set the clipping parameter $\epsilon$ to 0.15 to balance exploration and exploitation. In MSRA, the number of future steps is $h=3$ and the time-decay coefficient is $w=0.8$. In CORP, the self-reflection threshold is $\theta=0.06$. All other hyperparameters of these modules have been carefully tuned through sensitivity analysis, as detailed in Section $4.$
4.2. Experiment Performance (RQ1)
Table 2 summarizes the overall performance of our REG4Rec against multiple baselines on both public and industrial datasets. The best results are highlighted in bold, while the second-best results are underlined. From these results, we make the following observations:
@@ 4.2. Experiment Performance (RQ1)
4.3. Ablation Study (RQ2)
To assess the effectiveness of each component, we conduct ablation studies on the industrial dataset, which is the most complex and representative scenario on our recommendation platform. As shown in Table 1, the industrial dataset is orders of magnitude larger than public datasets, providing a more realistic and challenging evaluation environment.
  • w/o MPQ: Replaces MPQ with the widely used RQ-VAE (Yang et al., 2025; Zhang et al., 2025b). To ensure a fair comparison, in RQ-VAE, both the number of codebooks and the size of each codebook are kept consistent with those in MPQ. Additionally, the same multimodal representations are used as input for both models. In RQ-VAE, since the token for the item is sequential, the codebook order is fixed at inference.
@@ 4.3. Ablation Study (RQ2)
Table 3 presents the performance of these ablation experiments. We summarize the main observations below:
@@
Table 3. Ablation Study of REG4Rec.
@@
$4.$ Sensitivity Analysis (RQ3)
@@
Figure 3. Sensitivity Analysis
$x$.
4.5. Reasoning Step Scaling (RQ4)
@@
Figure 4. The performance of REG4Rec across different numbers of reasoning steps.
$x$.
4.6. Efficient Training Performance
We evaluate LADQ (Section  3.4.4) during training. Following prior work (Wang et al., 2018), we measure throughput as batch-samples per second (QPS) and report R@5 and N@5 for recommendation quality. As shown in Table 4, on the industrial dataset, LADQ reaches 4.5 QPS, representing a 32.3% speedup over the fp32 baseline. This acceleration comes at a negligible cost to accuracy, with only a 1.3% relative drop in R@5 and a 0.8% drop in N@5. In contrast, uniform quantization schemes (bf16 and fp8) offer higher QPS but incur larger accuracy degradation. Overall, LADQ offers a superior trade-off, achieving substantial training acceleration with minimal performance degradation.
Table 4. Training performance under different quantization schemes on the industrial dataset. The performance differences compared to the baseline (fp32) are highlighted in bold.
@@ 4.6. Efficient Training Performance
5. Online Experiments
To further validate our approach, we conducted an online A/B test on an advertising recommendation platform of a leading e-commerce company in Southeast Asia from July 18 to 22, 2025. The control group is based on SASRec, whereas the experimental group employs our proposed REG4Rec. Both groups consisted of 15% randomly selected users. Specifically, we observed 5.60% increase in the Advertising Revenue, 1.81% increase in the Click-Through Rate (CTR) and 3.29% increase in the Gross Merchandise Volume (GMV). The results of the online experiment once again confirm the performance of our proposed REG4Rec.
Table 5. Results of the online A/B experiment, with all performance gains being statistically significant at $p<0.05$.
@@ 5. Online Experiments
6. Conclusion and Future Work

### v4

@@ VI-B Inference with Reflection
Specifically, consistency is measured by JS divergence between consecutive distributions (Eq. (13)). If $JS(\tilde{p}_{i}\|\tilde{p}_{i+s})$ exceeds the self-reflection threshold $\theta$, it signals potential inconsistency or low confidence. The model then triggers self-reflection, terminating the current path and rolling back to an earlier step for regeneration. This pruning preserves only semantically coherent and logically consistent paths, enhancing overall recommendation accuracy.
VII Offline Experiments
@@
  • RQ5: How effective is our proposed training strategy in REG4Rec, LADQ, in balancing training efficiency and model accuracy?
  • RQ6: Can REG4Rec deliver performance improvements in real-world online recommendation platforms?
TABLE I: Statistics of Public and Industrial Datasets. #Users and #Items denote the number of users and items, respectively; #Interactions denotes the total number of user–item interactions in the dataset.
@@
VII-A Experimental Setup
Dataset. Four datasets are used to demonstrate the performance of our proposed method. The first three are public datasets, while the fourth one is an in-house industrial dataset. Table I summarizes dataset statistics.
  • Amazon Product Reviews. We use three publicly available benchmarks from the Amazon Product Reviews dataset [29], which contains user reviews and item meta information from May 1996 to July 2014. For GR, we focus on three categories: ”Beauty”, ”Sports and Outdoors”, and ”Toys and Games”. We generate training sequences from users’ review histories, sorted by timestamp, and exclude users with fewer than five reviews to ensure sufficient interactions. We follow the standard leave-one-out evaluation protocol [2]: the last item is used for testing, the second-to-last for validation, and all preceding items for training.
  • Industrial Dataset. We construct the in-house offline dataset from sequential user behaviors and feedback logs collected on the Alibaba’s e-commerce advertising platform. The dataset spans diverse recommendation scenarios, including homepage recommendations, product detail pages, and interactive gaming sections, capturing varied user behaviors. It contains data from 30 million users and 40 million advertisements collected between October 2024 and May 2025, offering a comprehensive view of real-world user behavior and ad content.
TABLE II: Performance Comparison across baselines. The best metric for each (dataset, metric) pair is highlighted in bold, while the second-best is underlined. The last column (Improv.) denotes the relative improvement of our REG4Rec over the best baseline.
Dataset Metric SASRec $S^{3}$-Rec TIGER COBRA ReaRec STREAM REG4Rec Improv.
Beauty R@5 0.0387 0.0387 0.0454 0.0537 0.0431 0.0538 0.0586 +8.92%
N@5 0.0249 0.0244 0.0321 0.0395 0.0276 0.0388 0.0416 +5.32%
R@10 0.0605 0.0647 0.0648 0.0725 0.0654 0.0691 0.0792 +9.24%
N@10 0.0318 0.0327 0.0384 0.0456 0.0375 0.0422 0.0493 +8.11%
Sports R@5 0.0233 0.0251 0.0264 0.0305 0.0256 0.0324 0.0372 +14.81%
N@5 0.0154 0.0161 0.0181 0.0215 0.0179 0.0232 0.0252 +8.62%
R@10 0.0350 0.0385 0.0400 0.0434 0.0393 0.0452 0.0527 +16.59%
N@10 0.0192 0.0204 0.0225 0.0257 0.0219 0.0271 0.0296 +9.23%
Toys R@5 0.0463 0.0443 0.0521 0.0619 0.0552 0.0644 0.0676 +4.97%
N@5 0.0306 0.0294 0.0371 0.0462 0.0374 0.0442 0.0493 +6.71%
R@10 0.0675 0.0700 0.0712 0.0781 0.0747 0.0799 0.0862 +7.88%
N@10 0.0374 0.0376 0.0432 0.0515 0.0435 0.0502 0.0574 +11.46%
Industrial Dataset R@5 0.0658 0.0759 0.0903 0.0963 0.0742 0.0981 0.1094 +11.52%
N@5 0.0453 0.0530 0.0594 0.0646 0.0517 0.0680 0.0738 +8.53%
R@10 0.0916 0.1066 0.1293 0.1397 0.0982 0.1422 0.1569 +10.34%
N@10 0.0564 0.0628 0.0715 0.0785 0.0603 0.0804 0.0890 +10.70%
@@ VII-A Experimental Setup
  • SASRec [2] : A transformer-based model that employs the self-attention mechanism to capture long-term dependencies.
  • $S^{3}$-Rec [27] : A transformer-based model that integrates self-supervised learning to enhance representation learning.
  • TIGER [18] : A GR model built on an encoder-decoder structure, leveraging semantic IDs (based on RQ-VAE) and autoregressive decoder to improve the generalizable of recommended items.
  • COBRA [10] : A GR model with a decoder-only structure, incorporating semantic IDs (based on RQ-VAE) and dense vectors to capture semantic insights and collaborative signals.
  • ReaRec [3] : A transformer-based model that extends SASRec to generate multiple implicit intermediate outcomes without explicitly defined inference paths.
  • STREAM [4] : A transformer-based model uses textually similar items as explicit reasoning steps, guided by the GRPO algorithm for reasoning.
Evaluation Metrics. For evaluation, we adopt the Recall@K (R@K) and NDCG@K (N@K) as standard metrics for recommendation, following [10], with $K$ set to 5 and 10. These metrics assess both the accuracy of item retrieval and the quality of the ranking.
Implementation Details. This section summarizes our implementation settings. The training process is executed on a distributed TensorFlow [30] platform, consisting of 10 parameter servers and 10 workers, each with 2 GPUs. To ensure a fair comparison with baselines, we configure the total number of transformer layers to 3, with one layer for the encoder and two layers for the decoder. The input dimension is 128 and the hidden size per transformer layer is 640. For MPQ, the hyperparameter $\alpha$ used in the training process of MPQ is set to 0.001. Because MPQ training is not the core focus and these weights have minimal impact once MPQ converges, we do not perform a sensitivity study on them. For flexible reasoning paths, MPQ uses 8 codebooks with 300 codewords each. The tokenizers process both text and image features for the industrial and public datasets. Text embeddings are generated using the pre-trained Qwen3-Embedding 7B model [31] and image embeddings are generated using the pre-trained Pailitao v8 model [43], both with a dimensionality of 256. We set the clipping parameter $\epsilon$ to 0.15 to balance exploration and exploitation. In MSRA, the number of future steps is $h=3$ and the time-decay coefficient is $w=0.8$. In CORP, the self-reflection threshold is $\theta=0.06$. All other hyperparameters of these modules have been carefully tuned through sensitivity analysis, as detailed in Section VII-D.
VII-B Experiment Performance (RQ1)
Table II summarizes the overall performance of our REG4Rec against multiple baselines on both public and industrial datasets. The best results are highlighted in bold, while the second-best results are underlined. From these results, we make the following observations:
@@ VII-B Experiment Performance (RQ1)
VII-C Ablation Study (RQ2)
To assess the effectiveness of each component, we conduct ablation studies on the industrial dataset, which is the most complex and representative scenario on our recommendation platform. As shown in Table I, the industrial dataset is orders of magnitude larger than public datasets, providing a more realistic and challenging evaluation environment.
  • w/o MPQ: Replaces MPQ with the widely used RQ-VAE [10, 4]. To ensure a fair comparison, in RQ-VAE, both the number of codebooks and the size of each codebook are kept consistent with those in MPQ. Additionally, the same multimodal representations are used as input for both models. In RQ-VAE, since the token for the item is sequential, the codebook order is fixed at inference.
@@ VII-C Ablation Study (RQ2)
Table III presents the performance of these ablation experiments. We summarize the main observations below:
@@
TABLE III: Ablation Study of REG4Rec.
@@
Figure 5: Sensitivity Analysis
VII-D Sensitivity Analysis (RQ3)
@@
VII-E Reasoning Step Scaling (RQ4)
@@
Figure 6: The performance of REG4Rec across different numbers of reasoning steps.
VII-F Efficient Training Performance (RQ5)
We evaluate LADQ (Section  VI-A) during training. Following prior work [40], we measure throughput as batch-samples per second (QPS) and report R@5 and N@5 for recommendation quality. As shown in Table IV, on the industrial dataset, LADQ reaches 4.5 QPS, representing a 32.3% speedup over the fp32 baseline. This acceleration comes at a negligible cost to accuracy, with only a 1.3% relative drop in R@5 and a 0.8% drop in N@5. In contrast, uniform quantization schemes (bf16 and fp8) offer higher QPS but incur larger accuracy degradation. Overall, LADQ offers a superior trade-off, achieving substantial training acceleration with minimal performance degradation.
TABLE IV: Training performance under different quantization schemes on the industrial dataset. The performance differences compared to the baseline (fp32) are highlighted in bold.
@@ VII-F Efficient Training Performance (RQ5)
VIII Online Experiments (RQ6)
To further validate our approach, we conducted an online A/B test on an advertising recommendation platform of Alibaba, a leading e-commerce company in Southeast Asia, from July 18 to 22, 2025. The control group is based on SASRec, whereas the experimental group employs our proposed REG4Rec. Both groups consisted of 15% randomly selected users. Specifically, we observed 5.60% increase in the Advertising Revenue, 1.81% increase in the Click-Through Rate (CTR) and 3.29% increase in the Gross Merchandise Volume (GMV). The results of the online experiment once again confirm the performance of our proposed REG4Rec.
TABLE V: Results of the online A/B experiment, with all performance gains being statistically significant at $p<0.05$.
@@ VIII Online Experiments (RQ6)
IX Conclusion and Future Work
@@
X Ethical Considerations
Throughout this research, we have strictly followed established ethical guidelines and best practices in data handling, experimentation, and system deployment. All experiments were conducted using widely available open-source datasets, as well as an anonymized industrial dataset, in which every piece of user-related information was thoroughly removed, masked, or aggregated to ensure full compliance with privacy regulations and to protect individual identities. Special care was taken to prevent any possibility of re-identification, including the removal of indirect identifiers and sensitive behavioral traces. From a security and safety perspective, we implemented checks to ensure that the recommendations generated by REG4Rec pose no risk to user well-being, digital safety, or platform integrity. We also verified that no outputs promote harmful, misleading, or unsafe content. These measures, together with our privacy-preserving practices, ensure REG4Rec contributes positively to both research and real-world applications.